# Moment bounds for condition numbers and singular values of high-dimensional Gaussian random matrices: Applications and limitations

### v1

@@ 1 Introduction
Given a data matrix $X_{n}\in\mathbb{R}^{n\times p}$, the associated Gram matrix is defined as $S_{n}:=\frac{1}{n}X_{n}^{\top}X_{n}$. Gram matrices arise naturally in regression, covariance estimation, and related statistical models, where $X_{n}$ serves as either a design matrix or an observation matrix. Asymptotic investigations of the corresponding classical statistical estimators typically assume that the rows of $X_{n}$ are independently sampled from a multivariate Gaussian distribution. Modern data applications frequently operate in regimes where the dimension $p$ is comparable to, or exceeds, the sample size $n$. Consequently, meaningful asymptotic investigations require frameworks in which $p$ grows with $n$. In such high-dimensional settings, statistical accuracy and computational stability depend critically on the spectral properties of $X_{n}$. In these settings, characterizing the expected prediction risk—an essential measure of estimation accuracy—frequently hinges on obtaining sharp moment bounds for the condition number, or for the extreme singular values of $X_{n}$ (see Section 3 for concrete examples). These considerations motivate the present study. We investigate arbitrary positive moments of the condition number and largest singular value of $X_{n}$, together with arbitrary negative moments of its smallest (non-zero) singular value, under regimes where $p/n\to\gamma\in[0,\infty)$. The condition number is a central object in numerical linear algebra, governing stability, sensitivity, and algorithmic convergence. Its probabilistic behavior has therefore received sustained attention in random matrix theory. Following the seminal work of Edelman [11], numerous studies have established several distributional properties and tail bounds for condition numbers of random matrices; see, for example, [12, 28, manriquemirón2023conditionnumberrandomtridiagonal, 19]. For matrices with independent identically distributed (i.i.d.) Gaussian entries, a particularly rich theory is available [12, 11, 9, 22]. However, to the best of our knowledge, existing results largely focus on tail probabilities or expectations of logarithmic functionals and do not directly yield conclusions regarding the behavior of general moments of the condition number.
@@
As our first contribution, we investigate this question within the framework of a random design matrix $X_{n}\in\mathbb{R}^{n\times p}$ whose rows are i.i.d. multivariate Gaussian. We impose standard bounded-eigenvalue assumptions on the covariance matrix of the underlying Gaussian distribution. For any fixed positive $r$, our results (see Theorem 1 and Lemma 2) show that when $p/n\to\gamma\in[0,\infty)\setminus\{1\}$ as $n\to\infty$, the $r^{\text{th}}$ moment of the condition number of $X_{n}$ remains uniformly bounded by a constant depending only on $r$ (and not on $n$). Moreover, Lemma 2 demonstrates that the expected condition number diverges as $\gamma\to 1$, revealing a phase-transition phenomenon consistent with—and extending—the double descent behavior anticipated in [23]. Although our analysis incorporates tail bounds derived in [9], the proof requires a sequence of additional, nontrivial arguments beyond those available in the existing literature.
@@
Next, in addition to the behavior of condition number moments arising in several statistical applications (see Section 3), the study of negative moments of the smallest singular value of a Gaussian random matrix plays a fundamental role in problems such as Gram matrix estimation (see, e.g., [30, 24, 26]). A classical result of [26] shows that for $n\times p$ matrices with independent sub-Gaussian entries, the smallest singular value is typically of the order of the rectangular spectral gap, namely $\sqrt{n}-\sqrt{p}$ when $n>p$, with high probability. More precisely, this associated small-ball probability bound consists of two components: an exponentially decaying term and an anti-concentration (small-ball) term reflecting the local mass behavior of the entry distribution. While this formulation applies broadly to sub-Gaussian ensembles, it is not sufficiently sharp in the Gaussian case for establishing bounds on negative moments, due to the presence of the exponentially small term (see Section 4.1). To address this limitation, we derive a explicit non-asymptotic bounds for negative moments of the smallest singular value of a Gaussian random matrix. Using this bound, we show that for any fixed positive $r$, the $r^{\text{th}}$ moment of $\frac{\sqrt{n}}{s_{\min}(X_{n})}$ (where $s_{\min}$ denotes the smallest non-zero eigenvalue) remains uniformly bounded by a constant depending only on $r$ (and not on $n$) under the regime $p/n\to\gamma\in[0,\infty)\setminus\{1\}$. Our analysis is based on a small-ball probability bound tailored to Gaussian matrices $X_{n}$ (see Theorem 3). Additionally, using a concentration inequality from [33], we establish a complementary and technically more direct result for positive moments of the largest singular value (see Lemma 4) under the same asymptotic regime $p/n\to\gamma\in[0,\infty)$.
@@
Section 3 presents several applications of the preceding results in statistical settings involving Gram matrices. In particular, in Section 3.1 we derive an explicit upper bound for the mean prediction risk (under Frobenius norm) of a ridge estimator in a general multi-response linear regression model with $p$ predictors and $q$ responses (Lemma 5). As a consequence, Theorem 6 establishes that this prediction risk converges to zero under a random design framework and the standard scaling condition $pq=o(n)$. As a second application, in Section 3.2 we consider the sample variance–covariance matrix, a canonical example of a Gram-type matrix. Theorem 7 shows that the moment bounds for the estimation error of its inverse exhibit the same non-asymptotic rate as those of the sample covariance matrix ([15]) itself under a Gaussian design. Finally, Section 3.3 investigates the computational implications of our moment bounds through Gram-type linear systems. We analyze the expected iteration complexity of gradient descent for solving systems involving the sample Gram matrix. Theorem 9 establishes that, under Gaussian designs, the expected iteration complexity remains uniformly bounded whenever $p/n\to\gamma\in[0,\infty)\setminus\{1\}$. Moreover, we show that the expected iteration complexity is fundamentally governed by moments of the condition number, thereby providing a direct link between spectral stability and computational performance. In contrast, as $\gamma\to 1$, the expected complexity diverges (Lemma 10), revealing a phase-transition phenomenon consistent with the double descent behavior of the condition number.
@@
In Section 4, we discuss possible extensions of our results beyond the Gaussian design and clarify the inherent limitations. While many high-dimensional bounds derived under Gaussianity extend to sub-Gaussian designs via standard tail arguments, the behavior of the inverse of the smallest singular value is considerably more delicate. In particular, although the moment bounds for the largest singular value in Lemma 4 extend to general sub-Gaussian settings (see Remark 2.7), we show that the inverse-moment and condition-number results of Theorem 3 and Theorem 1 do not hold in full generality for sub-Gaussian designs. Specifically, we construct an explicit continuous sub-Gaussian counterexample (see Proposition 11) for which all negative moments of the smallest (non-zero) singular value and all positive moments of the condition number of the corresponding design matrix are infinite. We further show that, under this construction, the associated inverse sample covariance estimation error also diverges. We relate these phenomena to the well-known role of small-ball probabilities governing the lower tail behavior of the smallest singular value [26], and explain why Gaussianity permits substantially sharper inverse-moment control than what is available under general sub-Gaussian small-ball theory.
@@
The remainder of the paper is organized as follows. Section 1.1 introduces the notation and basic definitions used throughout the paper. Bounds for the smallest singular value, largest singular value, and condition number are developed in Section 2, where we establish the principal moment results (Theorem 1 and Theorem 3) of the paper. Statistical implications of these bounds are examined in Section 3, including applications to prediction risk (Theorem 6), covariance estimation (Theorem 7), and computational complexity (Theorem 9). Extensions beyond the Gaussian design, together with a discussion of fundamental limitations and counterexamples, are provided in Section 4. Proofs of some of the auxiliary results are deferred to the supplementary material.
@@ 1.1 Notation and definitions
In the following notation, $I_{p}$ denotes the identity matrix of order $p$. For any matrix $A\in\mathbb{R}^{n\times p}$, let $s_{\max}(A)$ denote its largest singular value. If $n>p$ and $\mathrm{rank}(A)=p$, let $s_{\min}(A)$ denote the smallest singular value of $A$, and define the condition number by $\kappa(A)=\frac{s_{\max}(A)}{s_{\min}(A)}$. If $n\leq p$ and $\mathrm{rank}(A)=n$, let $s_{\min}(A)$ denote the smallest nonzero singular value of $A$ (equivalently, the $n$-th singular value), and then the condition number is again defined by $\kappa(A)=\frac{s_{\max}(A)}{s_{\min}(A)}$. If $A$ is a symmetric square matrix, then $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ denote the smallest and largest eigenvalues of $A$. Also, we define the range (or column space) of $A$ as
@@
We define the vectorization of the matrix $A$ as $\mbox{vec}(A)=(A_{11},\dots,A_{p1},A_{12},\dots,A_{pp})^{T}$.
@@ 2 Main results
In this section, we derive several new results establishing upper bounds for the moments of the condition number, as well as the minimum and maximum singular values, of a Gaussian random matrix in a growing-dimension asymptotic regime under a general covariance structure. Specifically, we consider a random matrix $\tilde{X}_{n}\in\mathbb{R}^{n\times p}$ whose rows are independent and follow a multivariate Gaussian distribution with covariance matrix $\tilde{\Sigma}_{n}$. The sequence of matrices $\{\tilde{\Sigma}_{n}\}_{n\geq 1}$ are assumed to satisfy
@@
for some finite universal constants $c_{m}$ and $c_{M}$. The regularity condition in (1) ensures that the sequence $\{\tilde{\Sigma}_{n}\}_{n\geq 1}$ remains uniformly well-conditioned [7], in the sense that its eigenvalues are bounded away from both 0 and $\infty$ uniformly over $n$. Assumptions of this type are standard in high-dimensional asymptotic analysis; see, for example, [7, 13, 36, 29]. Under this framework, we first establish Theorem 1, which demonstrates that the moments of the condition number $\kappa(\tilde{X}_{n})$ remain uniformly bounded over $n$, provided that $p/n\to\gamma\in[0,1)$.
@@
Before providing the formal proof, we present a high-level overview of the proof strategy. As noted in the introduction, [9] established an asymptotic tail bound for the condition number $\kappa(Z_{n})$ of an $n\times p$ Gaussian random matrix $Z_{n}$ with i.i.d. $\mathcal{N}(0,1)$ entries. In particular, for $t>p$, they showed that the tail probability $\mathbb{P}(\kappa(Z_{n})>t)$ exhibits polynomial decay of order $t^{-(n-p+1)}$. Although this result does not directly imply bounds of the form stated in Theorem 1, it provides a crucial starting point. When combined with additional nontrivial probabilistic arguments, it can be leveraged to derive moment bounds of a comparable nature for the condition number. This is formalised in the proof below.
@@
for some positive constants $c_{1}$ and $c_{2}$ (depending only on $\gamma$), where the last two inequalities follow from [33, Corollary 5.35] and $b_{n}^{2}=(\sqrt{n}-\sqrt{p})^{2}\sim(1-\sqrt{\gamma})^{2}n$ under $p/n\to\gamma\in[0,1)$.
Let $a_{n}:=\sqrt{n}+\sqrt{p}$. By the standard upper tail bound for $s_{\max}(Z_{n})$ ( see [33, Corollary 5.35]), there exist positive constants $c_{3}$ and $c_{4}$ (depending only on $\gamma$) and a constant $t_{\star}=t_{\star}(\gamma)\geq 1$ such that, for all sufficiently large $n$ and all $t\geq t_{\star}$, it holds that
@@
where $(t\sigma_{n}-a_{n})_{+}=\max\{(t\sigma_{n}-a_{n}),0\}$, and the last inequality follows from [33, Corollary 5.35] together with $a_{n}^{2}=(\sqrt{n}+\sqrt{p})^{2}\sim(1+\sqrt{\gamma})^{2}n$ under $p/n\to\gamma\in[0,1)$. Combining the above bounds yields that, for all $t\geq t_{\star}$ and all sufficiently large $n$,
@@
First note that (1) implies $\tilde{\Sigma}_{n}\succ 0$ for all $n$, and hence $\mathrm{rank}(\tilde{X}_{n})=\min(n,p)$ almost surely. In particular, in the wide-matrix regime $p>n$, the smallest nonzero singular value of $\tilde{X}_{n}$ equals the $n$-th singular value.
@@
    When $p/n\to\gamma\in(1,\infty)$, we have $n/p\to\gamma^{-1}\in(0,1)$. Since the nonzero singular values of $\tilde{X}_{n}$ and $\tilde{X}_{n}^{\top}$ coincide, it follows that $\kappa(\tilde{X}_{n})=\kappa(\tilde{X}_{n}^{\top})$. Therefore, applying Theorem 1 to the transposed matrix $\tilde{X}_{n}^{\top}$ yields the stated uniform bound.
@@
Uniform boundedness of the expected condition number alone is not sufficient for asymptotic evaluations in many statistical applications (see Section 3). In addition to the uniform boundedness of the condition number established in Theorem 1, such evaluations repeatedly rely on more refined spectral controls. Specifically, the developments in Section 3 require: (i) bounds on the negative moments of the smallest singular value , which govern the stability of inverses and lower spectral regularity, and (ii) bounds on the positive moments of the largest singular value, which control the magnitude of the spectral norm.
We begin with the lower end of the singular-value spectrum. Theorem 3 controls negative moments of $s_{\min}(X_{n})$ when $X_{n}$ has i.i.d. $\operatorname{\mathcal{N}}(0,1)$ entries. The bound has two parts: a leading term of order $(n-p-1)^{-r/2}$ and an additional term that decays exponentially fast in (npr+1)(n-p-r+1). The leading term captures the dominant dependence on the gap between sample size and dimension, while the exponentially small remainder reflects that extremely small values of $s_{\min}(X_{n})$ become rapidly less likely as npn-p grows. In particular, under the growing dimension regime i.e. when $p/n\to\gamma\in[0,1)$, the lemma implies a uniform bound on $\operatorname{\mathbb{E}}\big[(\sqrt{n}/s_{\min}(X_{n}))^{r}\big]$, which is the form needed later when inverse-type quantities appear.
@@
Theorem 3 extends directly to the regime $p/n\to\gamma\in(1,\infty)$. The proof proceeds similarly as in part (i) of Lemma 2 upon interchanging the roles of $n$ and $p$, with $s_{\min}(X_{n})$ replaced by the smallest nonzero singular value. Furthermore, invoking part (ii) of Lemma 2, one obtains that when $p/n\to 1$,
@@
We next control the upper end of the singular-value spectrum. Lemma 4 bounds positive moments of $s_{\max}(X_{n})$ when $X_{n}$ has i.i.d. $\operatorname{\mathcal{N}}(0,1)$ entries. It is well known that $s_{\max}(X_{n})$ typically grows on the order of $\sqrt{n}+\sqrt{p}$. Lemma 4 formalizes in the context of moment convergence this by showing that, for any fixed $r>0$, the $r$-th moment of $s_{\max}(X_{n})$ is controlled at the corresponding order, namely $\operatorname{\mathbb{E}}[s_{\max}(X_{n})^{r}]\lesssim(\sqrt{n}+\sqrt{p})^{r}$ (up to constants depending only on $r$). As a consequence, under the proportional growth regime $p/n\to\gamma\in[0,\infty)$, the normalized quantity $s_{\max}(X_{n})/\sqrt{n}$ has uniformly bounded moments.
@@
Now, by [33, Corollary 5.35], there exists an appropriate choice of $C$ (independent of $n$) such that
@@
Lemma 4 extends beyond the Gaussian setting. In particular, an analogous bound remains valid when the entries of $X_{n}$ are sub-Gaussian. This follows from standard nonasymptotic results for sub-Gaussian random matrices; see, for example, [33, Corollary 5.35] and its general formulation in [33, Theorem 5.39]. The primary modification required in the proof concerns the constants ( $c$ and $C$), which now depend on the sub-Gaussian norm of the underlying distribution (see Section 4). In contrast, Theorem 3 is essentially Gaussian-specific and does not, in general, extend to arbitrary sub-Gaussian designs. We refer to Section 4 for further discussion and an explicit counterexample.

### v2

# arXiv:2602.21487v2[math.ST] 27 Feb 2026

@@ 1 Introduction
Given a data matrix $X_{n}\in\mathbb{R}^{n\times p}$, the associated Gram matrix is defined as $S_{n}:=\frac{1}{n}X_{n}^{\top}X_{n}$  [^1] It is more common to refer to $nS_{n}=X_{n}^{\top}X_{n}$ as the Gram matrix, but we refer to $S_{n}$ as the Gram matrix for expositional convenience. Gram matrices arise naturally in regression, covariance estimation, and related statistical models, where $X_{n}$ serves as either a design matrix or an observation matrix. Asymptotic investigations of the corresponding classical statistical estimators typically assume that the rows of $X_{n}$ are independently sampled from a multivariate Gaussian distribution. Modern data applications frequently operate in regimes where the dimension $p$ is comparable to, or exceeds, the sample size $n$. Consequently, meaningful asymptotic investigations require frameworks in which $p$ grows with $n$. In such high-dimensional settings, statistical accuracy and computational stability depend critically on the spectral properties of $X_{n}$. In these settings, characterizing the expected prediction risk—an essential measure of estimation accuracy—frequently hinges on obtaining sharp moment bounds for the condition number, or for the extreme singular values of $X_{n}$ (see Section 3 for concrete examples). These considerations motivate the present study. We investigate arbitrary positive moments of the condition number and largest singular value of $X_{n}$, together with arbitrary negative moments of its smallest singular value, under regimes where $p/n\to\gamma\in[0,\infty)$.
The condition number is a central object in numerical linear algebra, governing stability, sensitivity, and algorithmic convergence. Its probabilistic behavior has therefore received sustained attention in random matrix theory. Following the seminal work of Edelman [11], numerous studies have established several distributional properties and tail bounds for condition numbers of random matrices; see, for example, [12, 27, manriquemirón2023conditionnumberrandomtridiagonal, 19]. For matrices with independent identically distributed (i.i.d.) Gaussian entries, a particularly rich theory is available [12, 11, 9, 22]. However, to the best of our knowledge, existing results largely focus on tail probabilities or expectations of logarithmic functionals and do not directly yield conclusions regarding the behavior of general moments of the condition number.
@@
As our first contribution, we investigate this question within the framework of a random design matrix $X_{n}\in\mathbb{R}^{n\times p}$ whose rows are i.i.d. multivariate Gaussian. We impose standard bounded-eigenvalue assumptions on the covariance matrix of the underlying Gaussian distribution. For any fixed positive $r$, our results (see Theorem 1 and Lemma 2) show that when $p/n\to\gamma\in[0,\infty)\setminus\{1\}$ as $n\to\infty$, the $r^{\text{th}}$ moment of the condition number of $X_{n}$ remains uniformly bounded by a constant depending only on $r$ (and not on $n$). Moreover, Lemma 2 demonstrates that the expected condition number diverges when $\gamma=1$, revealing a phase-transition phenomenon consistent with—and extending—the double descent behavior anticipated in [23]. Although our analysis incorporates tail bounds derived in [9], the proof requires a sequence of additional, nontrivial arguments beyond those available in the existing literature.
@@
Uniform boundedness of the expected condition number alone is not sufficient for asymptotic evaluations in many statistical applications. In particular, the study and control of negative moments of the smallest singular value of a Gaussian random matrix plays a fundamental role in problems such as Gram matrix estimation (see, e.g., [5, 10, 14, 24, 25]). A classical result of [25] shows that for $n\times p$ matrices with independent sub-Gaussian entries, the smallest singular value is typically of the order of the rectangular spectral gap, namely $\sqrt{n}-\sqrt{p}$ when $n>p$, with high probability. More precisely, the associated small-ball probability bound consists of two components: an exponentially decaying term and an anti-concentration (small-ball) term reflecting the local mass behavior of the entry distribution. While this formulation applies broadly to sub-Gaussian ensembles, it is not sufficiently sharp in the Gaussian case for establishing bounds on negative moments, due to the presence of the exponentially small term (see Section 4.1). To address this limitation, we derive explicit non-asymptotic bounds for negative moments of the smallest singular value of a Gaussian random matrix. Using these bounds, we show that for any fixed positive $r$, the $r^{\text{th}}$ moment of $\frac{\sqrt{n}}{s_{\min}(X_{n})}$ (where $s_{\min}$ denotes the smallest singular value) remains uniformly bounded by a constant depending only on $r$ (and not on $n$) under the regime $p/n\to\gamma\in[0,\infty)\setminus\{1\}$. Our analysis is based on a small-ball probability bound tailored to Gaussian matrices $X_{n}$ (see Theorem 3). Additionally, using a concentration inequality from [31], we establish a complementary and technically more direct result for positive moments of the largest singular value (see Lemma 4) under the same asymptotic regime $p/n\to\gamma\in[0,\infty)$.
@@
Section 3 presents several applications of the preceding results in statistical settings involving Gram matrices. In particular, in Section 3.1 we derive an explicit upper bound for the mean prediction risk (under Frobenius norm) of a ridge estimator in a general multi-response linear regression model with $p$ predictors and $q$ responses (Lemma 5). As a consequence, Theorem 6 establishes that this prediction risk converges to zero under a random design framework and the standard scaling condition $pq=o(n)$. As a second application, in Section 3.2 we consider the sample variance–covariance matrix, a canonical example of a Gram-type matrix. Theorem 7 shows that the moment bounds for the estimation error of its inverse exhibit the same non-asymptotic rate as those of the sample covariance matrix ([15]) itself under a Gaussian design. Finally, Section 3.3 investigates the computational implications of our moment bounds through Gram-type linear systems. We analyze the expected iteration complexity of gradient descent for solving systems involving the sample Gram matrix. Theorem 9 establishes that, under Gaussian designs, the expected iteration complexity remains uniformly bounded whenever $p/n\to\gamma\in[0,\infty)\setminus\{1\}$. Moreover, we show that the expected iteration complexity is fundamentally governed by moments of the condition number, thereby providing a direct link between spectral stability and computational performance. In contrast, as $p/n\to 1$, the expected complexity diverges (Lemma 10), revealing a phase-transition phenomenon consistent with the double descent behavior of the condition number.
@@
In Section 4, we discuss possible extensions of our results beyond the Gaussian design and clarify the inherent limitations. While many high-dimensional bounds derived under Gaussianity extend to sub-Gaussian designs via standard tail arguments, the behavior of the inverse of the smallest singular value is considerably more delicate. In particular, although the moment bounds for the largest singular value in Lemma 4 extend to general sub-Gaussian settings (see Remark 2.7), we show that the inverse-moment and condition-number results of Theorem 3 and Theorem 1 do not hold in full generality for sub-Gaussian designs. Specifically, we construct an explicit continuous sub-Gaussian counterexample (see Proposition 11) for which all negative moments of the smallest singular value and all positive moments of the condition number of the corresponding design matrix are infinite. We further show that, under this construction, the associated inverse sample covariance estimation error also diverges. We relate these phenomena to the well-known role of small-ball probabilities governing the lower tail behavior of the smallest singular value [25], and explain why Gaussianity permits substantially sharper inverse-moment control than what is available under general sub-Gaussian small-ball theory.
@@
The remainder of the paper is organized as follows. Section 1.1 introduces the notation and basic definitions used throughout the paper. The moment bounds for the smallest singular value and condition number (Theorems 3 and 1) are developed in Section 2. Statistical implications of these bounds are examined in Section 3, including applications to high-dimensional regression (Theorem 6), covariance estimation (Theorem 7), and computational complexity (Theorem 9). Extensions beyond the Gaussian design, together with a discussion of fundamental limitations and counterexamples, are provided in Section 4. Proofs of some of the auxiliary results are deferred to the supplementary material.
@@ 1.1 Notation and definitions
In the following notation, $I_{p}$ denotes the identity matrix of order $p$. For any matrix $A\in\mathbb{R}^{n\times p}$, let $s_{\max}(A)$ denote its largest singular value, and $s_{\min}(A)$ denote its $\min(n,p)^{th}$ singular value (when singular values are arranged in descending order). When $\mathrm{rank}(A)=\min(n,p)$, then $s_{\min}(A)$ corresponds to the smallest non-zero singular value of $A$. The condition number of $A$ is defined as $\kappa(A):=\frac{s_{\max}(A)}{s_{\min}(A)}$. If $A$ is a symmetric square matrix, then $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ denote the smallest and largest eigenvalues of $A$. Also, we define the range (or column space) of $A$ as
@@
We define the vectorization of the matrix $A$ as $\mbox{vec}(A)=(A_{11},\dots,A_{p1},A_{12},\dots,A_{pp})^{\top}$.
@@ 2 Main results
In this section, we derive several new results establishing upper bounds for the moments of the condition number, as well as the negative moments of the minimum singular value, of a Gaussian random matrix. These results are derived in a growing-dimension asymptotic regime under a general covariance structure. Specifically, we consider a random matrix $\tilde{X}_{n}\in\mathbb{R}^{n\times p}$ whose rows are independent and follow a multivariate Gaussian distribution with covariance matrix $\tilde{\Sigma}_{n}$. The sequence of matrices $\{\tilde{\Sigma}_{n}\}_{n\geq 1}$ are assumed to satisfy
@@
for some finite universal constants $c_{m}$ and $c_{M}$. The regularity condition in (1) ensures that the sequence $\{\tilde{\Sigma}_{n}\}_{n\geq 1}$ remains uniformly well-conditioned [7], in the sense that its eigenvalues are bounded away from both 0 and $\infty$ uniformly over $n$. Assumptions of this type are standard in high-dimensional asymptotic analysis; see, for example, [7, 13, 34, 28]. Note that, since the rows of $\tilde{X}_{n}$ are assumed to be independent Gaussian, under the assumption (1) we have $\operatorname{rank}(\tilde{X}_{n})=\min(n,p)$ almost surely. Hence, $s_{\min}(\tilde{X}_{n})$ denotes the smallest non-zero singular value of $\tilde{X}_{n}$ in this context. Under this framework, we first establish Theorem 1, which demonstrates that the moments of the condition number $\kappa(\tilde{X}_{n})$ remain uniformly bounded over $n$, provided that $p/n\to\gamma\in[0,1)$.
@@
Before providing the formal proof, we present a high-level overview of the proof strategy. As noted in the introduction, [9] established an asymptotic tail bound for the condition number $\kappa(Z_{n})$ of an $n\times p$ Gaussian random matrix $Z_{n}$ with i.i.d. $\mathcal{N}(0,1)$ entries. In particular, for $t>p$, they showed that the tail probability $\mathbb{P}(\kappa(Z_{n})>t)$ exhibits polynomial decay of order $t^{-(n-p+1)}$. Although this result does not directly imply bounds of the form stated in Theorem 1, it provides a crucial starting point. When combined with additional nontrivial probabilistic arguments, it can be leveraged to derive the required moment bounds for the condition number. This is formalized in the proof below.
@@
for some positive constants $c_{1}$ and $c_{2}$ (depending only on $\gamma$), where the last two inequalities follow from [31, Corollary 5.35] and $b_{n}^{2}=(\sqrt{n}-\sqrt{p})^{2}\sim(1-\sqrt{\gamma})^{2}n$ under $p/n\to\gamma\in[0,1)$.
Let $a_{n}:=\sqrt{n}+\sqrt{p}$. By the standard upper tail bound for $s_{\max}(Z_{n})$ ( see [31, Corollary 5.35]), there exist positive constants $c_{3}$ and $c_{4}$ (depending only on $\gamma$) and a constant $t_{\star}=t_{\star}(\gamma)\geq 1$ such that, for all sufficiently large $n$ and all $t\geq t_{\star}$, it holds that
@@
where $(t\sigma_{n}-a_{n})_{+}=\max\{(t\sigma_{n}-a_{n}),0\}$, and the last inequality follows from [31, Corollary 5.35] together with $a_{n}^{2}=(\sqrt{n}+\sqrt{p})^{2}\sim(1+\sqrt{\gamma})^{2}n$ under $p/n\to\gamma\in[0,1)$. Combining the above bounds yields that, for all $t\geq t_{\star}$ and all sufficiently large $n$,
@@
First recall that (1) implies $\tilde{\Sigma}_{n}\succ 0$ for all $n$, and hence $\mathrm{rank}(\tilde{X}_{n})=\min(n,p)$ almost surely.
@@
    When $p/n\to\gamma\in(1,\infty)$, we have $n/p\to\gamma^{-1}\in(0,1)$. Since the singular values of $\tilde{X}_{n}$ and $\tilde{X}_{n}^{\top}$ coincide, it follows that $\kappa(\tilde{X}_{n})=\kappa(\tilde{X}_{n}^{\top})$. Therefore, applying Theorem 1 to the transposed matrix $\tilde{X}_{n}^{\top}$ yields the stated uniform bound.
@@
Uniform boundedness of the condition number moments alone is not sufficient for asymptotic evaluations in many statistical applications (see Section 3). In addition to the uniform boundedness of the condition number established in Theorem 1, such evaluations repeatedly rely on more refined spectral controls. Specifically, the developments in Section 3 require: (i) bounds on the negative moments of the smallest singular value, which govern the stability of inverses and lower spectral regularity, and (ii) bounds on the positive moments of the largest singular value, which control the magnitude of the spectral norm.
We begin with the lower end of the singular-value spectrum. Theorem 3 controls negative moments of $s_{\min}(X_{n})$ when $X_{n}$ has i.i.d. $\operatorname{\mathcal{N}}(0,1)$ entries. The bound has two parts: a leading term of order $(n-p-1)^{-r/2}$ and an additional term that decays exponentially fast in (npr+1)(n-p-r+1). The leading term captures the dependence of the negative moments of the minimum singular value on the gap between the sample size and the dimension, while the exponentially small remainder reflects that extremely small values of $s_{\min}(X_{n})$ become rapidly less likely as npn-p grows. In particular, under the growing dimension regime i.e. when $p/n\to\gamma\in[0,1)$, the lemma implies a uniform bound on $\operatorname{\mathbb{E}}\big[(\sqrt{n}/s_{\min}(X_{n}))^{r}\big]$, which is the form needed later when inverse-type quantities appear.
@@
Theorem 3 extends directly to the regime $p/n\to\gamma\in(1,\infty)$. The proof proceeds similarly as in part (i) of Lemma 2 upon interchanging the roles of $n$ and $p$, and replacing $X_{n}$ by $X_{n}^{\top}$. Note that $X_{n}$ and $X_{n}^{\top}$ share the same singular values. Furthermore, invoking part (ii) of Lemma 2, one obtains that when $p/n\to 1$,
@@
We next control the upper end of the singular-value spectrum. Lemma 4 bounds positive moments of $s_{\max}(X_{n})$ when $X_{n}$ has i.i.d. $\operatorname{\mathcal{N}}(0,1)$ entries. It is well known that $s_{\max}(X_{n})$ typically grows on the order of $\sqrt{n}+\sqrt{p}$ ([31, 4]). Lemma 4 formalizes this in the context of moment convergence by showing that, for any fixed $r>0$, the $r$-th moment of $s_{\max}(X_{n})$ is controlled at the corresponding order, namely $\operatorname{\mathbb{E}}[s_{\max}(X_{n})^{r}]\lesssim(\sqrt{n}+\sqrt{p})^{r}$ (up to constants depending only on $r$). As a consequence, under the proportional growth regime $p/n\to\gamma\in[0,\infty)$, the normalized quantity $s_{\max}(X_{n})/\sqrt{n}$ has uniformly bounded moments.
@@
Now, by [31, Corollary 5.35], there exists an appropriate choice of $C$ (independent of $n$) such that
@@
Lemma 4 extends beyond the Gaussian setting. In particular, an analogous bound remains valid when the entries of $X_{n}$ are sub-Gaussian. This follows from standard nonasymptotic results for sub-Gaussian random matrices; see, for example, [31, Corollary 5.35] and its general formulation in [31, Theorem 5.39]. The primary modification required in the proof concerns the constants ($c$ and $C$), which now depend on the sub-Gaussian norm of the underlying distribution (see Section 4). In contrast, Theorem 1 and Theorem 3 are essentially Gaussian-specific and does not, in general, extend to arbitrary sub-Gaussian designs. We refer the reader to Section 4 for further discussion and an explicit counterexample.